### v16

# Rational angle bisectors on the coordinate plane and solutions of Pell’s equations

Abstract
On the coordinate plane, the slopes of two straight lines and the slope of one of their angle bisectors satisfy the equation Recently, an explicit formula for non-trivial integral solutions of this equation with solutions of negative Pell’s equations was discovered by the author. In this article, for a given square-free integer and a given integer we describe every integral solution of such that are coprime by using the fundamental unit of and elements of whose absolute value of norms are the smallest prime powers. We also describe every non-trivial rational solution of the above equation as one of its applications.
1 Introduction
On the coordinate plane, the following problem can be considered.
Problem 1.
For which rational numbers are the slopes of the angle bisectors between two straight lines with slopes rational?
Remark 1.
Given two straight lines, we consider the two angles formed by them, regardless of whether they are acute or not. The bisector of one of the angles and that of the supplementary angle are perpendicular to each other. In the case when they are not parallel to the coordinate axes, if one of the slopes is rational, then so is the other, since the product of the slopes is
Problem 1 has the meaning when the bisectors of can be drawn by connecting and other lattice points for given lattice points which is important in drawing techniques. Furthermore, in engineering, we can specify the radiation range and the axis of light with a ratio of integers without error due to approximations to irrational numbers by using a solution to this problem. Problem 1 is reduced to solving the equation
in (see [4, Proposition 1]). We say that a solution of ( ‣ 1) is trivial if
Example 1.
![[Uncaptioned image]](x1.png)
The following proposition plays an important role in solving Problem 1.
Proposition 1.
Every non-trivial rational solution of ( ‣ 1) is given by
for some rational solutions of
where is a square-free integer
The proof is given in Section 3. The following formula for integral solutions of ( ‣ 1) was proven by using certain criteria for the divisibility of integral solutions of Pell’s equations.
Theorem 1 ([4, Theorem 1]).
Throughout this article, we use the following definition and symbols.
Definition 1.
Let be a square-free integer and let be an integer
Let be an integral solution of We say that is strictly primitive if are coprime.
For each equation we call its positive integral solution with minimum its fundamental solution.
Remark 2.
Suppose that is a square number, and let be an integral solution of If is primitive, that is, are coprime, then is strictly primitive; otherwise, there exists a common prime divisor of and dividing twice and also twice, which contradicts that are coprime and is square-free.
Notation 1.
Let denote the additive monoid of all integers
Let be a square-free integer Let be the fundamental unit of the real quadratic field Let be the fundamental solution of and put
Put
For each put
and let be the fundamental solution of
and put
Put
Furthermore, for each where we denote the conjugate and norm of by
respectively.
The first main theorem of this article is the following formula.
Theorem 2.
Let be a non-trivial rational solution of ( ‣ 1).
Case : are the -components of rational solutions of the equation Then is given by
for some such that
Case : are the -components of rational solutions of a common negative Pell’s equation for some square-free integer Then is given by
for some such that and In addition, are written of the form
with some satisfying
and
The proof is given in Section 3.
Example 2.
The result in Case (I) refines the previous result [4, Theorem 2]. For any rational solution of ( ‣ 1) in Case (I), each of and is a ratio of the leg lengths of some Pythagorean triangle. We find the rational solutions of ( ‣ 1) in Case (II) by using the following Theorem. This is the second main theorem of this article.
Theorem 3.
Let be an integer
Suppose that or “ and ” or Then has a strictly primitive integral solution if and only if
(1.5) for some for each prime number where
Suppose that and Then has a strictly primitive integral solution if and only if
(1.6) for some for each prime number
In these cases, its strictly primitive integral solution satisfies
(1.7) for some such that
(1.8) and where in the latter case.
- (2)
The proof is given in Section 2.
2 Rational solutions of Pell’s equations
In this section, we prove Theorem 3. Throughout the section, let be a square-free integer and use the symbols in Notation 1. We recall the following classical theorem.
Theorem 4.
The positive Pell’s equation has a non-trivial integral solution independently of the value of
The negative Pell’s equation has a rational solution, if and only if has no prime factors congruent to modulo
Remark 3.
If has an integral solution, then has no prime factors congruent to modulo however, the converse is false. For example, has a rational solution but no integral solutions.
Let If have integral solutions respectively, then has an integral solution
since the norm map is a group homomorphism. In the case when such a solution is called a Pell multiple.
Theorem 5 ([3, Corollary 3.5]).
Let If has an integral solution, then there exists a finite number of its integral solutions such that its every integral solution satisfies
for some and
In this article, we generalize the concept of Pell multiples. The following propositions are fundamental.
Proposition 2.
Let be distinct prime numbers, and let If has a strictly primitive integral solution for each then defined by
(2.1) is a strictly primitive integral solution of
Let be a prime number, and let be integers If has a strictly primitive integral solution then defined by
(2.2) is a strictly primitive integral solution of
Proof.
We prove the contrapositions.
For each let be an integral solution of and put Suppose that the integral solution of defined by (2.1) is not strictly primitive. Then for some
Case: Then and therefore
Case: Then and therefore divides
This implies and therefore
Thus and therefore the solution is not strictly primitive.
Let be an integral solution of Suppose that the solution of defined by (2.2) is not strictly primitive. Then
Case: Then and therefore
Case: Then and therefore With respect to an extension of the -adic additive valuation to this implies
Hence
Thus and therefore the solution is not strictly primitive.∎
Corollary 1.
Let Suppose that has an integral solution. If has a strictly primitive integral solution, then so does
Proof.
If has a strictly primitive integral solution then defined by is a strictly primitive integral solution of ∎
We say that an ideal in or the integer ring of a quadratic field is primitive if is not divisible by for any prime number
Theorem 6.
Let be a prime number. Suppose that or Then has a strictly primitive integral solution for some multiple of if and only if splits in
Suppose that
If then
Both and hold if and only if In this case, holds.
If and is a multiple of then every integral solution of is a pair of even numbers, and not strictly primitive.
If or “ and ” or then
If and then
Unless “ and and ” then is not greater than the ideal class number of for each
Proof.
Put Recall that the integer ring of is
Let be the monoid of all ideals in and let be the equivalence relation on defined by
which is compatible with the multiplication. Since the ideal class group of is finite, and so is the quotient monoid which can be regarded as its submonoid. Furthermore, splits in if and is unramified in if
Suppose that a prime number splits into prime ideals and in that is, Then there exist such that
by the finiteness of where are coprime since is primitive. Furthermore, are coprime, since and These imply that has a strictly primitive integral solution.
To prove the converse, suppose that an integer has a prime factor which does not split in Let be an integral solution of
Case: is unramified in or “ and is unramified in ” In the principal ideal is decomposed as
where both sides are divisible by Therefore and are divisible by in since the prime ideal divides one of these ideals and also the other because of its self-conjugacy. This implies that and are divisible by in since
(2.3) by the assumption. Hence and therefore
Case: ramifies in Then and therefore divides which implies
In these cases, is not strictly primitive. Thus the assertion of (1) holds.
Suppose that Then has no integral solutions, since
This implies that if
Suppose that Then the pair of odd integers defined by satisfies This implies that are coprime, and therefore and In this case, holds by the minimality of
Conversely, if and then a strictly primitive integral solution of satisfies
which implies since
If then
which imply that every integral solution of is a pair of even numbers for each multiple of
If and then since has no strictly primitive integral solutions for each integer by (2). Combining this fact with (1) and (2), we obtain the desired result. ∎
Theorem 7.
Let be pairwise coprime ideals in the integer ring of a quadratic field Then is primitive if and only if are primitive.
Proof.
For any ideal in is primitive if and only if the residue group is cyclic (see [1, Corollary 6.30]). Combining this fact with the Chinese remainder theorem, we obtain the desired assertion. ∎
Now we are ready to prove Theorem 3.
Proof of Theorem 3.
Put and let denote the integer ring of For each prime number which splits in we take a prime ideal in lying over in and put Note that are primitive for such a prime ideal in and any Let be a strictly primitive integral solution of Let be the set of all prime factors of and put For each let and be the quotient and remainder, respectively, when dividing by For each put The equality can be expressed as with ideals in which implies
since for each For each if or then is decomposed as
in by Theorem 6 (1), and therefore the ideal is divisible by either or ; otherwise, in and also by (2.3), and therefore which contradicts that are coprime. For each if or we can see that is divisible by either or by the same argument, which implies
- (i)
Case: and Then is unramified in and by Theorem 6. The prime ideal divides both and only once by its self-conjugacy; otherwise, in and therefore in since
(2.4) which contradicts that are coprime. The ideal is decomposed as
with in The ideal
is principal and primitive by Theorem 7, and therefore equal to which implies that for each Hence there exists such that
Letting we obtain (1.7).
- (iii)
Case: and Then and has no strictly primitive integral solutions for each multiple of by Theorem 6, which implies The ideal is decomposed as
with in The ideal
is principal and primitive by Theorem 7, and therefore equal to which implies that for each Hence there exists such that
Letting
and defining by
where the first double sign is not necessarily the same as the second, we obtain (1.7).
The exponent satisfies (1.8), since for each and
by (see [6, Theorem 2.1.4]). Thus the assertion of (1) holds.
Let be an integral solution of and let be the greatest common divisor of and Let be the set of all prime factors of and put . Put
Then is a strictly primitive integral solution of and satisfies
(2.5) for some satisfying (1.5), (1.6), (1.8) by (1). Put
Case or . Then and since
Case . Then and since and
Case or . Then
Case . Then
Multiplying both sides of (2.5) by
we obtain (1.9).∎
Remark 4.
By Theorem 3, it is impossible for both and to have strictly primitive integral solutions, if has no integral solutions.
Example 3.
The equation has the fundamental solution however, has no integral solutions. The fundamental unit of is The equations have the fundamental solutions respectively. Every integral solution of satisfies one of the following conditions for some
Theorem 3 implies the following formula for rational solutions of Pell’s equations.
Theorem 8.
Let Every rational solution of satisfies
for some such that
and
Proof.
Every integral solution of satisfies
for some
Let be a rational solution of where are coprime integers and Then is an integral solution of and satisfies
for some where
and
by Theorem 3. Dividing both sides by we obtain (2.6). Condition (2.7) follows from
and
We can remove Condition (2.8), since and for each Thus the assertion of the theorem holds. ∎
3 Rational angle bisectors
Proof of Proposition 1.
Let be a non-trivial rational solution of ( ‣ 1), and let
with Substituting these into ( ‣ 1) and multiplying both sides by we obtain
The combinations of prime factors of which divide odd times coincide with each other, since are square numbers. Let be the product of these prime factors (define if there are no such prime factors). Then there exist such that
where are the -components of the integral solutions of and are the -components of the rational solutions of with Substituting (3.2) into (3.1) and dividing both sides by we obtain
or equivalently
Solving for we obtain
since and therefore
Proof of Theorem 2.
Case : are the -components of rational solutions of respectively. Let
with Then we have
Furthermore, put
Then we have
These imply
and therefore
Case : are the -components of rational solutions of for some square-free integer respectively. In general, if and then
Combining these identities with Theorem 8, we obtain the desired assertion.∎
4 Appendix
In Theorem 2, for such that has a rational solution and the values of are summarized in the following table. For reference, we add the values of the ideal class number the fundamental unit of and below the values of respectively.

### v17

@@
 Abstract
-On the coordinate plane, the slopes of two straight lines and the slope of one of their angle bisectors satisfy the equation Recently, an explicit formula for non-trivial integral solutions of this equation with solutions of negative Pell’s equations was discovered by the author.
+On the coordinate plane, the slopes of two straight lines and the slope of one of their angle bisectors satisfy the equation Recently, an explicit formula for nontrivial integral solutions of this equation with solutions of negative Pell’s equations was discovered by the author.
 In this article, for a given square-free integer and a given integer we describe every integral solution of such that are coprime by using the fundamental unit of and elements of whose absolute value of norms are the smallest prime powers.
-We also describe every non-trivial rational solution of the above equation as one of its applications.
+We also describe every nontrivial rational solution of the above equation as one of its applications.
 1 Introduction
-On the coordinate plane, the following problem can be considered.
+On the coordinate plane, we consider the following problem, which can be called the rational angle bisection problem.
 Problem 1.
@@
 Problem 1 is reduced to solving the equation
-in (see [4, Proposition 1]). We say that a solution of ( ‣ 1) is trivial if
+in (see [5, Proposition 1]). We say that a solution of ( ‣ 1) is trivial if
 Example 1.
-![[Uncaptioned image]](x1.png)
 The following proposition plays an important role in solving Problem 1.
 Proposition 1.
-Every non-trivial rational solution of ( ‣ 1) is given by
+Every nontrivial rational solution of ( ‣ 1) is given by
 for some rational solutions of
@@
 The proof is given in Section 3. The following formula for integral solutions of ( ‣ 1) was proven by using certain criteria for the divisibility of integral solutions of Pell’s equations.
-Theorem 1 ([4, Theorem 1]).
+Theorem 1 ([5, Theorem 1]).
 Throughout this article, we use the following definition and symbols.
@@
 Let be an integral solution of We say that is strictly primitive if are coprime.
-For each equation we call its positive integral solution with minimum its fundamental solution.
+Let be positive integral solutions of We say is smaller than if
+For each equation we call its minimum positive integral solution its fundamental solution.
 Remark 2.
 Suppose that is a square number, and let be an integral solution of If is primitive, that is, are coprime, then is strictly primitive; otherwise, there exists a common prime divisor of and dividing twice and also twice, which contradicts that are coprime and is square-free.
+In particular, every integral solution of is strictly primitive.
 Notation 1.
@@
 Theorem 2.
-Let be a non-trivial rational solution of ( ‣ 1).
+Let be a nontrivial rational solution of ( ‣ 1).
 Case : are the -components of rational solutions of the equation Then is given by
@@
 Example 2.
-The result in Case (I) refines the previous result [4, Theorem 2].
+The result in Case (I) refines the previous result [5, Theorem 2].
 For any rational solution of ( ‣ 1) in Case (I), each of and is a ratio of the leg lengths of some Pythagorean triangle.
@@
 Theorem 4.
-The positive Pell’s equation has a non-trivial integral solution independently of the value of
+The positive Pell’s equation has a nontrivial integral solution independently of the value of
 The negative Pell’s equation has a rational solution, if and only if has no prime factors congruent to modulo
+Proof.
+See [1, Theorem 3.2.1].
+By Fermat’s two squares theorem (see [8, Theorem 2.15]), can be expressed as for some if and only if has no prime factors congruent to modulo We can assume since Under these conditions, is a rational solution of
+Conversely, if has a rational solution with then which implies that has no prime factors congruent to modulo by Fermat’s theorem as above, and so does ∎
 Remark 3.
@@
 since the norm map is a group homomorphism. In the case when such a solution is called a Pell multiple.
-Theorem 5 ([3, Corollary 3.5]).
+Theorem 5.
 Let If has an integral solution, then there exists a finite number of its integral solutions such that its every integral solution satisfies
 for some and
+Proof.
+See [4, Corollary 3.5]. ∎
 In this article, we generalize the concept of Pell multiples. The following propositions are fundamental.
@@
 Case: Then and therefore
-Case: Then and therefore divides
-This implies and therefore
+Case: Then This implies that divides
+and therefore in Hence
 Thus and therefore the solution is not strictly primitive.
@@
 Case: Then and therefore
-Case: Then and therefore With respect to an extension of the -adic additive valuation to this implies
+Case: Then and therefore in With respect to an extension of the -adic additive valuation to this implies
 Hence
 Thus and therefore the solution is not strictly primitive.∎
-Corollary 1.
+Proposition 3.
 Let Suppose that has an integral solution. If has a strictly primitive integral solution, then so does
 Proof.
-If has a strictly primitive integral solution then defined by is a strictly primitive integral solution of ∎
+We prove the contraposition.
+Suppose that has no strictly primitive integral solutions.
+Let be an integral solution of Then defined by is an integral solution of which implies that there exists a prime divisor of such that
+Case: Then and therefore
+Case: Then In this implies and therefore since Hence
+Thus This implies that has no strictly primitive integral solutions. ∎
 We say that an ideal in or the integer ring of a quadratic field is primitive if is not divisible by for any prime number
@@
 If and then
-Unless “ and and ” then is not greater than the ideal class number of for each
+Unless “ and ” then
+for each
 Proof.
@@
 which imply that every integral solution of is a pair of even numbers for each multiple of
-If and then since has no strictly primitive integral solutions for each integer by (2). Combining this fact with (1) and (2), we obtain the desired result. ∎
+If and then since has no strictly primitive integral solutions for each integer by (2). Combining this fact with (1) and (2), we obtain the desired equalities.
+The desired inequality follows from the argument in the proof of (1).∎
 Theorem 7.
@@
 Proof.
-For any ideal in is primitive if and only if the residue group is cyclic (see [1, Corollary 6.30]). Combining this fact with the Chinese remainder theorem, we obtain the desired assertion. ∎
+For any ideal in is primitive if and only if the residue group is cyclic (see [2, Corollary 6.30]). Combining this fact with the Chinese remainder theorem, we obtain the desired assertion. ∎
 Now we are ready to prove Theorem 3.
@@
 The exponent satisfies (1.8), since for each and
-by (see [6, Theorem 2.1.4]). Thus the assertion of (1) holds.
+by (see [7, Theorem 2.1.4]). Thus the assertion of (1) holds.
 Let be an integral solution of and let be the greatest common divisor of and Let be the set of all prime factors of and put . Put
 Then is a strictly primitive integral solution of and satisfies
-(2.5) for some satisfying (1.5), (1.6), (1.8) by (1). Put
+(2.5) for some satisfying (1.8) and the conditions obtained by replacing with in (1.5), (1.6) because of (1). Put
 Case or . Then and since
@@
 Proof of Proposition 1.
-Let be a non-trivial rational solution of ( ‣ 1), and let
+Let be a nontrivial rational solution of ( ‣ 1), and let
 with Substituting these into ( ‣ 1) and multiplying both sides by we obtain
@@
 Combining these identities with Theorem 8, we obtain the desired assertion.∎
-4 Appendix
+Appendix
 In Theorem 2, for such that has a rational solution and the values of are summarized in the following table.

### v18

@@
 Abstract
-On the coordinate plane, the slopes of two straight lines and the slope of one of their angle bisectors satisfy the equation Recently, an explicit formula for nontrivial integral solutions of this equation with solutions of negative Pell’s equations was discovered by the author.
-In this article, for a given square-free integer and a given integer we describe every integral solution of such that are coprime by using the fundamental unit of and elements of whose absolute value of norms are the smallest prime powers.
+On the coordinate plane, the slopes and of two straight lines and the slope of one of their angle bisectors satisfy the equation Recently, an explicit formula for nontrivial integral solutions of this equation with solutions of negative Pell’s equations was discovered by the author.
+In this article, for a given square-free integer and a given integer we describe every integral solution of such that and are coprime by using the fundamental unit of and elements of whose absolute value of norms are the smallest prime powers.
 We also describe every nontrivial rational solution of the above equation as one of its applications.
@@
 In the case when they are not parallel to the coordinate axes, if one of the slopes is rational, then so is the other, since the product of the slopes is
-Problem 1 has the meaning when the bisectors of can be drawn by connecting and other lattice points for given lattice points which is important in drawing techniques.
-Furthermore, in engineering, we can specify the radiation range and the axis of light with a ratio of integers without error due to approximations to irrational numbers by using a solution to this problem.
+Essentially, Problem 1 has the meaning when the bisectors of can be drawn by connecting and other lattice points for given lattice points which is important in drawing techniques.
+Furthermore, in engineering, we can specify the radiation range and the axis of light with a ratio of integers without errors due to approximations to irrational numbers by using a solution to this problem.
 Problem 1 is reduced to solving the equation
@@
 Example 1.
+![[Uncaptioned image]](x1.png)
 The following proposition plays an important role in solving Problem 1.
@@
 for some rational solutions of
-where is a square-free integer
+where is a positive square-free integer.
 The proof is given in Section 3. The following formula for integral solutions of ( ‣ 1) was proven by using certain criteria for the divisibility of integral solutions of Pell’s equations.
 Theorem 1 ([5, Theorem 1]).
+For each square-free integer such that has an integral solution, we denote the -th smallest positive integral solution of by Every nontrivial integral solution of ( ‣ 1) is given by
+for some integers after switching and if necessary, where (1.1) contains the case when Conversely, every triple of Form (1.1) or (1.2) is an integral solution of ( ‣ 1).
 Throughout this article, we use the following definition and symbols.
 Definition 1.
-Let be a square-free integer and let be an integer
+Let be a square-free integer, and let be an integer.
 Let be an integral solution of We say that is strictly primitive if are coprime.
-Let be positive integral solutions of We say is smaller than if
+Let be positive integral solutions of We say is smaller than if or “ and ”
 For each equation we call its minimum positive integral solution its fundamental solution.
@@
 Notation 1.
-Let denote the additive monoid of all integers
-Let be a square-free integer Let be the fundamental unit of the real quadratic field Let be the fundamental solution of and put
-Put
-For each put
+Let denote the additive monoid of all nonnegative integers. For each prime number let be the normalized -adic additive valuation. We denote the greatest common divisor of by
+Notation 2.
+Let be a square-free integer. Let be the fundamental unit of the real quadratic field Let be the fundamental solution of and let Let
+For each let
 and let be the fundamental solution of
-and put
-Put
-Furthermore, for each where we denote the conjugate and norm of by
+and let
+Let
+Furthermore, for each where we denote the conjugate, norm of by
 respectively.
@@
 Let be a nontrivial rational solution of ( ‣ 1).
-Case : are the -components of rational solutions of the equation Then is given by
-for some such that
-Case : are the -components of rational solutions of a common negative Pell’s equation for some square-free integer Then is given by
-for some such that and In addition, are written of the form
+Suppose that are the -components of rational solutions of the equation Then is given by
+(1.3) for some such that
+Suppose that are the -components of rational solutions of a common negative Pell’s equation for some square-free integer Then is given by
+(1.4) for some such that and In addition, can be written in the form
 with some satisfying
 and
+Conversely, every triple of Form (1.3) or (1.4) is a rational solution of ( ‣ 1).
 The proof is given in Section 3.
 Example 2.
-The result in Case (I) refines the previous result [5, Theorem 2].
-For any rational solution of ( ‣ 1) in Case (I), each of and is a ratio of the leg lengths of some Pythagorean triangle.
-We find the rational solutions of ( ‣ 1) in Case (II) by using the following Theorem.
+The result in (i) refines the previous result [5, Theorem 2].
+For any rational solution of ( ‣ 1) in (i), each of and is a ratio of the leg lengths of some Pythagorean triangle.
+We find the rational solutions of ( ‣ 1) in (ii) by using the following Theorem.
 This is the second main theorem of this article.
 Theorem 3.
-Let be an integer
-Suppose that or “ and ” or Then has a strictly primitive integral solution if and only if
-(1.5) for some for each prime number where
-Suppose that and Then has a strictly primitive integral solution if and only if
-(1.6) for some for each prime number
-In these cases, its strictly primitive integral solution satisfies
+Let be an integer. Then has a strictly primitive integral solution if and only if
+(1.5) for some for each prime number where “ and ” if In this case, its strictly primitive integral solution satisfies
 (1.6) for some such that
-(1.8) and where in the latter case.
-- (2)
+(1.7) and
+Let be an integer. Then every integral solution of satisfies
+(1.8) for some where (1.7) and
+(1.9) hold.
 The proof is given in Section 2.
 2 Rational solutions of Pell’s equations
-In this section, we prove Theorem 3. Throughout the section, let be a square-free integer and use the symbols in Notation 1. We recall the following classical theorem.
+In this section, we prove Theorem 3. Throughout the section, let be a square-free integer. We recall the following classical theorems.
+Theorem 4 ([1, Theorem 3.2.1]).
+The positive Pell’s equation has a nontrivial integral solution independently of the value of
 Theorem 5.
-The positive Pell’s equation has a nontrivial integral solution independently of the value of
-The negative Pell’s equation has a rational solution, if and only if has no prime factors congruent to modulo
+The negative Pell’s equation has a rational solution, if and only if has no prime divisors congruent to modulo
 Proof.
-See [1, Theorem 3.2.1].
-By Fermat’s two squares theorem (see [8, Theorem 2.15]), can be expressed as for some if and only if has no prime factors congruent to modulo We can assume since Under these conditions, is a rational solution of
-Conversely, if has a rational solution with then which implies that has no prime factors congruent to modulo by Fermat’s theorem as above, and so does ∎
+By Fermat’s two squares theorem (see [8, Theorem 2.15]), can be expressed as for some if and only if has no prime divisors congruent to modulo We can assume since Under these conditions, is a rational solution of
+Conversely, if has a rational solution with then which implies that has no prime divisors congruent to modulo by Fermat’s theorem as above, and so does ∎
 Remark 3.
-If has an integral solution, then has no prime factors congruent to modulo however, the converse is false. For example, has a rational solution but no integral solutions.
+If has an integral solution, then has no prime divisors congruent to modulo however, the converse is false. For example, has a rational solution but no integral solutions.
 Let If have integral solutions respectively, then has an integral solution
 since the norm map is a group homomorphism. In the case when such a solution is called a Pell multiple.
-Theorem 5.
+Theorem 6 ([4, Corollary 3.5]).
 Let If has an integral solution, then there exists a finite number of its integral solutions such that its every integral solution satisfies
 for some and
-Proof.
-See [4, Corollary 3.5]. ∎
 In this article, we generalize the concept of Pell multiples. The following propositions are fundamental.
@@
 (2.1) is a strictly primitive integral solution of
-Let be a prime number, and let be integers If has a strictly primitive integral solution then defined by
+Let be a prime number, and let be integers. If has a strictly primitive integral solution then defined by
 (2.2) is a strictly primitive integral solution of
@@
 We prove the contrapositions.
-For each let be an integral solution of and put Suppose that the integral solution of defined by (2.1) is not strictly primitive. Then for some
-Case: Then and therefore
-Case: Then This implies that divides
-and therefore in Hence
+For each let be an integral solution of and let Suppose that the integral solution of defined by (2.1) is not strictly primitive. Then for some
+Case 1: Suppose that Then and therefore
+Case 2: Suppose that Then In this implies that divides
+and therefore Hence
 Thus and therefore the solution is not strictly primitive.
 Let be an integral solution of Suppose that the solution of defined by (2.2) is not strictly primitive. Then
-Case: Then and therefore
-Case: Then and therefore in With respect to an extension of the -adic additive valuation to this implies
+Case 1: Suppose that Then and therefore
+Case 2: Suppose that Then and therefore in With respect to an extension of to this implies
 Hence
@@
 Let be an integral solution of Then defined by is an integral solution of which implies that there exists a prime divisor of such that
-Case: Then and therefore
-Case: Then In this implies and therefore since Hence
+Case 1: Suppose that Then and therefore
+Case 2: Suppose that Then In this implies and therefore since Hence
 Thus This implies that has no strictly primitive integral solutions. ∎
@@
 Let be a prime number. Suppose that or Then has a strictly primitive integral solution for some multiple of if and only if splits in
-Suppose that
-If then
-Both and hold if and only if In this case, holds.
-If and is a multiple of then every integral solution of is a pair of even numbers, and not strictly primitive.
-If or “ and ” or then
-If and then
-Unless “ and ” then
-for each
 Proof.
-Put Recall that the integer ring of is
-Let be the monoid of all ideals in and let be the equivalence relation on defined by
-which is compatible with the multiplication.
+Let Recall that the integer ring of is
+Let be the monoid of all ideals in For each we define if
+for some Then the relation on is an equivalence relation which is compatible with the multiplication.
 Since the ideal class group of is finite, and so is the quotient monoid which can be regarded as its submonoid.
@@
 by the finiteness of where are coprime since is primitive. Furthermore, are coprime, since and These imply that has a strictly primitive integral solution.
-To prove the converse, suppose that an integer has a prime factor which does not split in Let be an integral solution of
-Case: is unramified in or “ and is unramified in ” In the principal ideal is decomposed as
+To prove the converse, suppose that an integer has a prime divisor which does not split in Let be an integral solution of
+Case 1: Suppose that is unramified in Then or “ and ” by assumption. In the principal ideal is decomposed as
 where both sides are divisible by Therefore and are divisible by in since the prime ideal divides one of these ideals and also the other because of its self-conjugacy.
 This implies that and are divisible by in since
-(2.3) by the assumption. Hence and therefore
-Case: ramifies in Then and therefore divides which implies
-In these cases, is not strictly primitive. Thus the assertion of (1) holds.
-Suppose that Then has no integral solutions, since
-This implies that if
+(2.4) by assumption. Hence and therefore
+Case 2: Suppose that ramifies in Then and therefore divides which implies
+In these cases, is not strictly primitive. Thus the desired assertion holds. ∎
+Corollary 1.
+Unless “ and ” then is not greater than the ideal class number of for each
+Proof.
+The desired assertion follows from (2.3). ∎
+In the arguments below, we use the following lemma.
+Lemma 1.
+Suppose that and let be an integer.
+If then has no integral solutions. Furthermore, if then
+Both and hold if and only if In this case, holds.
+If and then every integral solution of is a pair of even integers. Furthermore, if then
+If and then every integral solution of is a pair of even integers.
+Proof.
+If then has no integral solutions, since by assumption. This implies that if
 Suppose that Then the pair of odd integers defined by satisfies This implies that are coprime, and therefore and In this case, holds by the minimality of
-Conversely, if and then a strictly primitive integral solution of satisfies
-which implies since
+Conversely, if and then a strictly primitive integral solution of satisfies which implies since
+(2.5) -
+Suppose that Then
+which imply that every integral solution of is a pair of even integers if This implies by (2).
 If then
-which imply that every integral solution of is a pair of even numbers for each multiple of
-If and then since has no strictly primitive integral solutions for each integer by (2). Combining this fact with (1) and (2), we obtain the desired equalities.
-The desired inequality follows from the argument in the proof of (1).∎
+which imply that every integral solution of is a pair of even integers if
+Now we can determine the elements of
+Theorem 8.
+If or “ and ” or then consists of all prime numbers which split in
+If and then consists of all prime numbers which split in and
+Proof.
+We also use the following fact.
 Theorem 9.
@@
 Proof of Theorem 3.
-Put and let denote the integer ring of For each prime number which splits in we take a prime ideal in lying over in and put Note that are primitive for such a prime ideal in and any Let be a strictly primitive integral solution of Let be the set of all prime factors of and put For each let and be the quotient and remainder, respectively, when dividing by For each put The equality can be expressed as with ideals in which implies
+Let and let denote the integer ring of For each prime number we take a prime ideal in lying over in and let Note that are primitive for each if splits in Let be a strictly primitive integral solution of Let be the set of all prime divisors of and let For each let and be the quotient and remainder, respectively, when dividing by For each let The equality can be expressed as with ideals in which implies
 since for each For each if or then is decomposed as
-in by Theorem 6 (1), and therefore the ideal is divisible by either or ; otherwise, in and also by (2.3), and therefore which contradicts that are coprime.
-For each if or we can see that is divisible by either or by the same argument, which implies
-- (i)
-Case: and Then is unramified in and by Theorem 6. The prime ideal divides both and only once by its self-conjugacy; otherwise, in and therefore in since
-(2.4) which contradicts that are coprime. The ideal is decomposed as
+in by Theorem 7, and therefore the ideal is divisible by either or ; otherwise, in and also by (2.4), and therefore which contradicts that are coprime.
+For each we can see that is divisible by either or by the same argument as above, which implies
+The latter case follows from since
+(2.6) and the fundamental solution of is strictly primitive.
+Case 2: Suppose that and
+Assume that Then is unramified in by Theorem 8, and by Lemma 1 (1), (4).
+The prime ideal divides both and only once by its self-conjugacy; otherwise, in and therefore in by (2.6), which contradicts that are coprime.
+The ideal is decomposed as
 with in The ideal
 is principal and primitive by Theorem 9, and therefore equal to which implies that for each Hence there exists such that
-Letting we obtain (1.7).
-- (iii)
-Case: and Then and has no strictly primitive integral solutions for each multiple of by Theorem 6, which implies The ideal is decomposed as
-with in The ideal
-is principal and primitive by Theorem 7, and therefore equal to which implies that for each Hence there exists such that
-Letting
-and defining by
-where the first double sign is not necessarily the same as the second, we obtain (1.7).
-The exponent satisfies (1.8), since for each and
-by (see [7, Theorem 2.1.4]). Thus the assertion of (1) holds.
-Let be an integral solution of and let be the greatest common divisor of and Let be the set of all prime factors of and put . Put
+This implies and therefore which contradicts that are coprime.
+Thus and therefore by Theorem 8.
+In these cases, the exponent satisfies (1.7), since
+by (2.5) and (see [7, Theorem 2.1.4]). We can verify that (1.5) is satisfied by the argument above. Thus the assertion of (1) holds.
+Let be an integral solution of and let Let be the set of all prime divisors of and let . Let
 Then is a strictly primitive integral solution of and satisfies
-(2.5) for some satisfying (1.8) and the conditions obtained by replacing with in (1.5), (1.6) because of (1). Put
-Case or . Then and since
-Case . Then and since and
-Case or . Then
-Case . Then
-Multiplying both sides of (2.5) by
+(2.7) for some satisfying (1.7) and the condition obtained by replacing with in (1.5) because of (1). If then (1.9) holds, since If then Multiplying both sides of (2.7) by
 we obtain (1.8).∎
@@
 with Substituting these into ( ‣ 1) and multiplying both sides by we obtain
-The combinations of prime factors of which divide odd times coincide with each other, since are square numbers.
-Let be the product of these prime factors (define if there are no such prime factors).
+For any prime number the parities of and coincide with each other, since are square numbers.
+Let be the product of every prime number such that these valuations are odd (define if there are no such prime numbers).
 Then there exist such that
@@
 Proof of Theorem 2.
-Case : are the -components of rational solutions of respectively. Let
+Suppose that are the -components of rational solutions of respectively. Let
 with Then we have
-Furthermore, put
+Furthermore, let
 Then we have
@@
 and therefore
-Case : are the -components of rational solutions of for some square-free integer respectively. In general, if and then
+Suppose that are the -components of rational solutions of for some square-free integer respectively. In general, if and then
 Combining these identities with Theorem 10, we obtain the desired assertion.∎

### v20

@@
 Problem 1.
-For which rational numbers are the slopes of the angle bisectors between two straight lines with slopes rational?
+For which rational numbers and are the slopes of the angle bisectors between two straight lines with slopes and rational?
 Remark 1.
@@
 In the case when they are not parallel to the coordinate axes, if one of the slopes is rational, then so is the other, since the product of the slopes is
-Essentially, Problem 1 has the meaning when the bisectors of can be drawn by connecting and other lattice points for given lattice points which is important in drawing techniques.
+Essentially, Problem 1 has the meaning when the bisectors of can be drawn by connecting and other lattice points for given lattice points and which is important in drawing techniques.
 Furthermore, in engineering, we can specify the radiation range and the axis of light with a ratio of integers without errors due to approximations to irrational numbers by using a solution to this problem.
@@
 Let be a square-free integer, and let be an integer.
-Let be an integral solution of We say that is strictly primitive if are coprime.
-Let be positive integral solutions of We say is smaller than if or “ and ”
-For each equation we call its minimum positive integral solution its fundamental solution.
+Let be an integral solution of We say that is strictly primitive, if and are coprime.
+Let be positive integral solutions of We say is smaller than if or if and
+For each equation of and we call its minimum positive integral solution its fundamental solution.
 Remark 2.
-Suppose that is a square number, and let be an integral solution of If is primitive, that is, are coprime, then is strictly primitive; otherwise, there exists a common prime divisor of and dividing twice and also twice, which contradicts that are coprime and is square-free.
+Suppose that is a square number, and let be an integral solution of If is primitive, that is, and are coprime, then is strictly primitive; otherwise, there exists a common prime divisor of and dividing twice and also twice, which contradicts that and are coprime, and is square-free.
 In particular, every integral solution of is strictly primitive.
@@
 Let
-Furthermore, for each where we denote the conjugate, norm of by
+Furthermore, for each where we denote the conjugate and norm of by
 respectively.
@@
 Let be a nontrivial rational solution of ( ‣ 1).
-Suppose that are the -components of rational solutions of the equation Then is given by
-(1.3) for some such that
-Suppose that are the -components of rational solutions of a common negative Pell’s equation for some square-free integer Then is given by
-(1.4) for some such that and In addition, can be written in the form
-with some satisfying
+Suppose that and are the -components of rational solutions of the equation Then is given by
+(1.3) for some such that and
+Suppose that and are the -components of rational solutions of a common negative Pell’s equation for some square-free integer Then is given by
+(1.4) for some such that and In addition, and can be written in the form
+for some and satisfying
 and
@@
 Let be an integer. Then has a strictly primitive integral solution if and only if
-(1.5) for some for each prime number where “ and ” if In this case, its strictly primitive integral solution satisfies
-(1.6) for some such that
-(1.7) and
+(1.5) for some for each prime number where and if In this case, its strictly primitive integral solution satisfies
+(1.6) for some and such that
+(1.7) -
 Let be an integer. Then every integral solution of satisfies
-(1.8) for some where (1.7) and
-(1.9) hold.
+(1.8) for some and where
+(1.9) and (1.7) hold.
 The proof is given in Section 2.
@@
 By Fermat’s two squares theorem (see [8, Theorem 2.15]), can be expressed as for some if and only if has no prime divisors congruent to modulo We can assume since Under these conditions, is a rational solution of
-Conversely, if has a rational solution with then which implies that has no prime divisors congruent to modulo by Fermat’s theorem as above, and so does ∎
+Conversely, if has a rational solution with and then which implies that has no prime divisors congruent to modulo by Fermat’s theorem as above, and so does ∎
 Remark 3.
 If has an integral solution, then has no prime divisors congruent to modulo however, the converse is false. For example, has a rational solution but no integral solutions.
-Let If have integral solutions respectively, then has an integral solution
+Let If and have integral solutions and respectively, then has an integral solution
 since the norm map is a group homomorphism. In the case when such a solution is called a Pell multiple.
 Theorem 6 ([4, Corollary 3.5]).
-Let If has an integral solution, then there exists a finite number of its integral solutions such that its every integral solution satisfies
+Let Suppose that has an integral solution. Then, for this equation, there exists a finite number of integral solutions such that every integral solution satisfies
 for some and
@@
 We prove the contrapositions.
-For each let be an integral solution of and let Suppose that the integral solution of defined by (2.1) is not strictly primitive. Then for some
+For each let be an integral solution of and let Suppose that the integral solution of defined by (2.1) is not strictly primitive. Then and for some
 Case 1: Suppose that Then and therefore
 Case 2: Suppose that Then In this implies that divides
-and therefore Hence
-Thus and therefore the solution is not strictly primitive.
-Let be an integral solution of Suppose that the solution of defined by (2.2) is not strictly primitive. Then
+and therefore Hence and
+Thus, in each case, and which imply that the solution is not strictly primitive.
+Let be an integral solution of Suppose that the solution of defined by (2.2) is not strictly primitive. Then and
 Case 1: Suppose that Then and therefore
 Case 2: Suppose that Then and therefore in With respect to an extension of to this implies
-Hence
-Thus and therefore the solution is not strictly primitive.∎
+Hence and
+Thus, in each case, and which imply that the solution is not strictly primitive.∎
 Proposition 3.
@@
 Suppose that has no strictly primitive integral solutions.
-Let be an integral solution of Then defined by is an integral solution of which implies that there exists a prime divisor of such that
+Let be an integral solution of Then defined by is an integral solution of which implies that there exists a prime divisor of such that and
 Case 1: Suppose that Then and therefore
-Case 2: Suppose that Then In this implies and therefore since Hence
-Thus This implies that has no strictly primitive integral solutions. ∎
+Case 2: Suppose that Then In this implies and therefore since Hence and
+Thus, in each case, and which imply that has no strictly primitive integral solutions. ∎
 We say that an ideal in or the integer ring of a quadratic field is primitive if is not divisible by for any prime number
@@
 Let Recall that the integer ring of is
-Let be the monoid of all ideals in For each we define if
+Let be the monoid of all nonzero ideals in For each we define if
 for some Then the relation on is an equivalence relation which is compatible with the multiplication.
@@
 Suppose that a prime number splits into prime ideals and in that is, Then there exist such that
-by the finiteness of where are coprime since is primitive. Furthermore, are coprime, since and These imply that has a strictly primitive integral solution.
+by the finiteness of where and are coprime since is primitive. Furthermore, and are coprime, since and These imply that has a strictly primitive integral solution.
 To prove the converse, suppose that an integer has a prime divisor which does not split in Let be an integral solution of
@@
 This implies that and are divisible by in since
-(2.4) by assumption. Hence and therefore
+(2.4) Hence and therefore and
 Case 2: Suppose that ramifies in Then and therefore divides which implies
-In these cases, is not strictly primitive. Thus the desired assertion holds. ∎
+Thus, in each case, is not strictly primitive. This proves the desired assertion. ∎
 Corollary 1.
-Unless “ and ” then is not greater than the ideal class number of for each
+Unless and then is not greater than the ideal class number of for each
 Proof.
@@
 If then has no integral solutions, since by assumption. This implies that if
-Suppose that Then the pair of odd integers defined by satisfies This implies that are coprime, and therefore and In this case, holds by the minimality of
+Suppose that Then the pair of odd integers defined by satisfies This implies that and are coprime, and therefore and In this case, holds by the minimality of
 Conversely, if and then a strictly primitive integral solution of satisfies which implies since
@@
 Theorem 8.
-If or “ and ” or then consists of all prime numbers which split in
+If or and or then consists of all prime numbers which split in
 If and then consists of all prime numbers which split in and
@@
 Proof of Theorem 3.
-Let and let denote the integer ring of For each prime number we take a prime ideal in lying over in and let Note that are primitive for each if splits in Let be a strictly primitive integral solution of Let be the set of all prime divisors of and let For each let and be the quotient and remainder, respectively, when dividing by For each let The equality can be expressed as with ideals in which implies
+Let and let denote the integer ring of For each prime number we take a prime ideal in lying over in and let Note that and are primitive for each if splits in Let be a strictly primitive integral solution of Let be the set of all prime divisors of and let For each let and be the quotient and remainder, respectively, when dividing by For each let The equality can be expressed as with ideals in which implies
 since for each For each if or then is decomposed as
-in by Theorem 7, and therefore the ideal is divisible by either or ; otherwise, in and also by (2.4), and therefore which contradicts that are coprime.
+in by Theorem 7, and therefore the ideal is divisible by either or ; otherwise, in and also by (2.4), and therefore and which contradict that and are coprime.
 For each we can see that is divisible by either or by the same argument as above, which implies
@@
 Case 2: Suppose that and
-Assume that Then is unramified in by Theorem 8, and by Lemma 1 (1), (4).
-The prime ideal divides both and only once by its self-conjugacy; otherwise, in and therefore in by (2.6), which contradicts that are coprime.
+Assume that Then is unramified in by Theorem 8, and by Lemma 1 (1) and (4).
+The prime ideal divides both and only once by its self-conjugacy; otherwise, in and therefore in by (2.6), which contradicts that and are coprime.
 The ideal is decomposed as
-with in The ideal
+with and in The ideal
 is principal and primitive by Theorem 9, and therefore equal to which implies that for each Hence there exists such that
-This implies and therefore which contradicts that are coprime.
+This implies and therefore and which contradict that and are coprime.
 Thus and therefore by Theorem 8.
+Subcase 2-2: Suppose that Then is unramified in by Lemma 1 (3), and and by Lemma 1 (2) and (4). The ideal is decomposed as
+with and in The ideal
+is principal and primitive by Theorem 9, and therefore equal to which implies that for each Hence there exists such that
+Letting for each and we obtain (1.6).
 In these cases, the exponent satisfies (1.7), since
@@
 Then is a strictly primitive integral solution of and satisfies
-(2.7) for some satisfying (1.7) and the condition obtained by replacing with in (1.5) because of (1). If then (1.9) holds, since If then Multiplying both sides of (2.7) by
+(2.7) for some and satisfying (1.7) and the condition obtained by replacing with in (1.5) because of (1). If then (1.9) holds, since If then Multiplying both sides of (2.7) by
 we obtain (1.8).∎
@@
 The equation has the fundamental solution however, has no integral solutions.
-The fundamental unit of is The equations have the fundamental solutions respectively.
+The fundamental unit of is The equations and have the fundamental solutions and respectively.
 Every integral solution of satisfies one of the following conditions for some
@@
 Let Every rational solution of satisfies
-for some such that
+for some and such that
 and
@@
 for some
-Let be a rational solution of where are coprime integers and Then is an integral solution of and satisfies
-for some where
+Let be a rational solution of where and are coprime integers and Then is an integral solution of and satisfies
+for some and where
 and
@@
 Let be a nontrivial rational solution of ( ‣ 1), and let
-with Substituting these into ( ‣ 1) and multiplying both sides by we obtain
-For any prime number the parities of and coincide with each other, since are square numbers.
+with and Substituting these into ( ‣ 1) and multiplying both sides by we obtain
+For any prime number the parities of and coincide with each other, since and are square numbers.
 Let be the product of every prime number such that these valuations are odd (define if there are no such prime numbers).
 Then there exist such that
-where are the -components of the integral solutions of and are the -components of the rational solutions of with Substituting (3.2) into (3.1) and dividing both sides by we obtain
-or equivalently
+where and are the -components of the integral solutions of Furthermore, and are the -components of the rational solutions of with and Substituting (3.2) into (3.1) and dividing both sides by we obtain
+or equivalently,
 Solving for we obtain
@@
 Proof of Theorem 2.
-Suppose that are the -components of rational solutions of respectively. Let
-with Then we have
+Suppose that and are the -components of rational solutions and of respectively. Let
+with and Then we have
 Furthermore, let
@@
 and therefore
-Suppose that are the -components of rational solutions of for some square-free integer respectively. In general, if and then
+which implies
+Suppose that and are the -components of rational solutions and of for some square-free integer respectively. In general, if and with then
+which imply
 Combining these identities with Theorem 10, we obtain the desired assertion.∎
 Appendix
-In Theorem 2, for such that has a rational solution and the values of are summarized in the following table.
-For reference, we add the values of the ideal class number the fundamental unit of and below the values of respectively.
+In Theorem 2, for such that has a rational solution and the values of and are summarized in the following table.
+For reference, we add the values of and below the values of and respectively.
+We also add the value of the ideal class number of